OpenStudy (babynini):

convert the equation to polar form x^2-y^2=7

OpenStudy (babynini):

(use variables r and theta as needed)

OpenStudy (babynini):

@jim_thompson5910 :)

OpenStudy (alekos):

substitute x=rcosθ and y=rsinθ

OpenStudy (babynini):

and that would be the final answer?

OpenStudy (babynini):

\[(rcos \theta)^2- (rsin \theta)^2=7\]

OpenStudy (babynini):

@alekos

OpenStudy (alekos):

No, I would go further \[r^{2}(\cos ^{2}\theta-\sin ^{2}\theta)=7\]

OpenStudy (alekos):

and the trig identity can now be simplified

OpenStudy (babynini):

\[r^2(\cos^2 \theta) = 7\]

OpenStudy (alekos):

no way. what is cos2θ ?

OpenStudy (babynini):

haha oops sorry I means cos2theta

OpenStudy (babynini):

\[r^2(\cos2 \theta) = 7\]

OpenStudy (alekos):

very good and now we can go further \[r ^{2}=7\sec2\theta \]

OpenStudy (babynini):

hmm I just need to write it in polar form. Should we be going this far?

OpenStudy (babynini):

did you mean cos?

OpenStudy (babynini):

oo I see what you did, okay, nevermind.

OpenStudy (alekos):

we are heading to the final polar form

OpenStudy (alekos):

just take the square root of both sides and you have the final polar form

OpenStudy (babynini):

\[r=\sqrt{7\sec2 \theta}\]

OpenStudy (alekos):

beautiful. thats it!

OpenStudy (babynini):

yay! thank you :)

OpenStudy (babynini):

hrmm they don't like that answer -.-

OpenStudy (babynini):

@Miracrown

OpenStudy (babynini):

oh, I figured it out :) they wanted r^2(cos2theta)=7